In a recent year, the Better Business Bureau settled 75% ofcomplaints they received. (Source: USA Today, March 2, 2009) You have been hired by the Bureau to investigate complaints this year involving computer stores. You plan to select a random sample of complaints to estimate the proportion of complaints the Bureau is able to settle. Assume the population proportion of complaints settled for the computer stores is the 0.75, as mentioned above. Suppose your sample size is 106. What is the probability that the sample proportion will be within 7 percent of the population proportion? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations Answer 00485(Enter your answer as a number accurate to 4 decimal places.)
Answer #1

here p =0.75

n = 106

P(sample proportion within 7% of population proportion )=

P(0.75-0.07 <p^ < 0.75+0.07) = P(0.68<p^ < 0.82)

standard deviation of sampling distribution of sample proportion = √(p(1-p)/n) = √(0.75*0.25/106) =0.042058

p =    0.75                              
std error,SE =    0.042058   
we need to calculate probability for ,                                  
0.68   ≤ p^ ≤    0.82                          
p1^ =    0.68   , p2^ =   0.82                  
                                  
Z1 =   (p1^ - p ) / SE =   -1.664                          
Z2 =   (p2^ - p ) / SE =   1.664                          
                                  
P (   0.68   < p^ <    0.82   ) =    P (    -1.664 < Z <    1.664   )
                                  
= P ( Z <    1.664   ) - P ( Z <   -1.664   ) =    0.9520   -    0.0480   =    0.9040(answer)